c)    If the area of a capacitor plates is squared three times and its square root distance tripled twice. What is the final value of its capacitance?   


1
Expert's answer
2021-09-23T18:00:34-0400

"C=\frac{A\epsilon_0}{d}"

"A'={3A}\\"

"d'=2d"

"C'=\frac{3A\epsilon_0}{2d}"

"\frac{C'}{C}=\frac{\frac{3A\epsilon_0}{2d}}{\frac{A\epsilon_0}{d}}=\frac{3}{2}"

"C'=\frac{3}{2}C"